OpenStudy (anonymous):

HELP FOR MDEAL PLZZ NEED HELP Purvi and her friends tipped their waiter a total of eight dollars for dinners that cost nine dollars , eight dollars , eleven dollars and twelve dollars each. Use a percent proportion to solve for the percent tip they left the waiter. Show your work

OpenStudy (anonymous):

9 + 8 + 11 + 12 = 40 (total of the meals) 8/40 = x/100 40x = 800 x = 800/40 x = 20%

OpenStudy (anonymous):

THANK YOU

OpenStudy (anonymous):

YW :)